# LLaNA: Large Language and NeRF Assistant

### v1

@@ 1 Introduction
Large Language Models (LLMs) [70, 1] have revolutionized the field of Natural Language Processing, demonstrating incredible text comprehension and generation capabilities. These results have fostered the development of Multimodal LLMs (MLLMs) [14, 81, 44, 11, 10], which can ingest various modalities such as images, videos, and audio, to generate then text describing and reasoning about the content of such modalities. Recently, MLLMs have also been extended to 3D data [58, 77], primarily represented through colored point clouds, yielding remarkable results even in this scenario.
@@
Figure 1: LLaNA. A new Multimodal Large Language Model that understands and reasons on an input NeRF. Notably, our framework processes directly the NeRF weights and performs tasks such as captioning, Q&A, and zero-shot classification of NeRFs.
@@ 2 Related work
Multi-modal Large Language Models (MLLMs).
@@
The interaction between NeRF and language has been recently investigated for several practical applications. Many works address the problem of generating geometrically consistent views of objects or scenes described by textual prompts [66, 49, 31, 65, 40, 36, 57]. Other approaches focus on editing the scene represented by a NeRF from text, e.g., by changing the appearance and shape of objects [73, 28, 67, 74, 69, 21, 79, 86], or by inserting/removing objects in the scene [3, 51]. Some techniques investigate new types of radiance fields that predict language features for each spatial location alongside density and color [32, 34]. In particular, by distilling knowledge from vision-language models into these models, the neural fields can be queried by textual prompts. Unlike all previous methods, Ballerini et al. [5] is the first to utilize NeRFs as an input modality. They aim to learn a mapping between the NeRF and CLIP [59] embedding spaces to perform tasks such as NeRF retrieval from textual or image queries. Differently, our goal is to develop an MLLM capable of reasoning about NeRFs.
@@ 3 Methodology
Neural Radiance Field (NeRF) [50] is a framework that employs coordinate-based neural networks, typically MultiLayer Perceptrons (MLP) and is trained on a collection of images of an object or scene taken from various vantage points. The main application of NeRFs is the task of novel views synthesis, i.e., photorealistic rendering of images from viewpoints unseen at training time. In its base formulation, the MLP is a function of continuous 3D coordinates $\mathbf{p}=(x,y,z)\in\mathbb{R}^{3}$, that yields four-dimensional outputs, $RGB\sigma\in[0,1]^{4}$. This output encodes the $RGB$ color and the volume density $\sigma$ of each 3D location in the scene. The volume density $\sigma$ can be interpreted as the differential probability of a ray terminating at point $\mathbf{p}$. After training, a NeRF can render images from any desired viewpoints at arbitrary resolution by querying it for the values of $RGB$ and $\sigma$ at several points along the ray corresponding to each pixel and applying the volumetric rendering equation [50].
In this work, we realize NeRFs as MLPs composed of $L$ hidden layers, an input layer, and an output layer. An example of MLP with 1 input, 1 output, and 1 hidden layer is shown in Fig. 2(left). A layer is parameterized by a weight matrix plus a bias vector. More in detail, the hidden layers in our architecture have the same number of input and output neurons, $H$, thus having squared weight matrices $\mathbf{W}_{l}\in\mathbb{R}^{H\times H}$ for $l=1,\dots,L$ and $H$-dimensional biases $\mathbf{b}_{l}\in{R}^{H}$. As input $\mathbf{p}$ goes through a 24-frequency encoding [50], the first layer has $\mathbf{W}_{in}\in\mathbb{R}^{144\times H}$ and $\mathbf{b}_{in}\in\mathbb{R}^{H}$. The final one has $\mathbf{W}_{out}\in\mathbb{R}^{H\times 4}$ and $\mathbf{b}_{out}\in\mathbb{R}^{4}$. Refer to Appendix A for more details on NeRFs.
@@
Inspired by the recent development of meta-networks capable of processing neural fields [42, 80], we employ as our meta-encoder architecture nf2vec [80]. It takes as input the weights of a NeRF and yields a global embedding that distills the content of the input NeRF. In particular, the weight matrices and biases of the input NeRF are stacked along the row dimension to form a matrix $\mathbf{M}\in\mathbb{R}^{S\times H}$, where $S=144+1+L*(H+1)+H+1=L*H+L+H+146$. Before stacking, we pad the output layer weights $\mathbf{W}_{out}$ and biases $\mathbf{b}_{out}$ with zeros to obtain $H$ columns (see Fig. 2, center).
@@
To train our framework, we hold multi-turn conversations about each NeRF available in the ShapeNeRF–Text dataset that we created (see Sec. 4). These conversations are organized into a set of prompts from the user and expected ground-truth answers that are used to optimize the original auto-regressive objective of the LLM. For the meta-encoder, we employ the nf2vec encoder pre-trained on ShapeNet released by the authors [80], and we keep it frozen during training. We follow the two-stage training protocol delineated in Liu et al. [44]:
@@
Our model is implemented in PyTorch and trained on 4 NVIDIA A100 with 64GB of VRAM each. Completing both stages requires $\sim$1 day of training. When fine-tuning LLaMA 2 for our specific applications, we strictly use a dataset that does not come from scraped web data. This approach should ensure that the integrity and effectiveness of the pre-existing safeguards are maintained.
@@ 4.1 ShapeNeRF–Text dataset
To train and validate our NeRF assistant, we automatically created a dataset of conversations about NeRFs, the ShapeNeRF–Text dataset. It features paired NeRFs and language annotations for ShapeNet objects [8], in particular for all the 40K NeRFs available in the nf2vec dataset [61].
Figure 3: Automatic annotation pipeline.
We followed the structure defined in PointLLM [77] to create the textual annotations. More in detail, we generated a brief description, a detailed description, 3 single-round Q&As, and one multi-round Q&A for each object. The brief descriptions are concise captions of the object, taking into account its global structure and appearance. The detailed descriptions are longer sentences that describe all the details of the object. The single-round Q&As consist of a question about the object and the corresponding ground-truth answer. Finally, the multi-round Q&As are longer conversations formed by 3 questions and the relative answers. The automatic data annotation pipeline is inspired by Cap3D [46] and is shown in Fig. 3. First, multiple views of each ShapeNet object have been rendered from different perspectives. Then, each view has been provided as input to LLaVA (LLaVA2-13b) [44] to get a detailed description of the object from that point of view. Afterward, starting from the captions generated by LLaVA, LLaMA 3 (LLaMA3-8B-chat) was used to generate the final ground-truth text data (brief and detailed descriptions, single and multi-round Q&As). Both the frozen LLMs employed to create our benchmark (LLaVA2-13b, LLaMA3-8b-chat) are equipped with safeguards.
When building the ground-truth data, to ensure diversity in the language annotations, each brief and detailed description has been associated with a question randomly sampled from 30 instructions for each kind of description. Such instructions, together with the carefully engineered request prompts for LLaVA and LLaMA, are reported in the supplementary material in Appendix C. ShapeNeRF–Text comprises 40K objects, with 240K textual annotations: 40K brief captions, 40K detailed descriptions, 120K single-round Q&A, and 40K three-round Q&A. The supplementary material shows an example of the resulting texts in Fig. 8. We split the dataset into train, validation, and test sets, stratified on the ShapeNet classes, containing distinct objects in proportions of 80%, 10%, and 10% of the total number of shapes, respectively.
@@ 4.2 Language tasks and metrics
We evaluate NeRF assistants on three different language tasks, given an input NeRF: brief captioning, detailed captioning, and single-round Q&A. We evaluate all tasks on the objects from the ShapeNeRF–Text test set. For brief captioning, we additionally evaluate the methods on the GPT2Shape Human Shape Text (HST) dataset [2], a subset of ShapeNet for which human-curated brief descriptions are publicly available. To generate the dialogues for HST, we randomly pair each of its shapes with one of the 30 instructions requesting a brief description. We employ standard language similarity metrics to evaluate these methods. We compute the cosine similarity between the global embeddings of the generated and ground-truth sentences provided by the pre-trained encoders Sentence-BERT [63] and SimCSE [18]. These metrics based on learned networks are the most effective at measuring the quality of the generated output. We also include standard handcrafted metrics based on n-gram statistics, like BLEU-1 [55], ROUGE-L [43], and METEOR [6].
@@ 5.1 Baselines
As our method is the first to investigate language tasks on NeRF, there are no baselines in the literature. However, given a NeRF, a straightforward way to create an assistant for it could be to render an image and use an MLLM capable of ingesting images. Alternatively, we could extract the 3D shape from the NeRF and use one of the recent 3D MLLMs. Hence, we use MLLMs as off-the-shelf foundation models, trained on hundreds of thousands of shapes or millions of images, without performing any fine-tuning on the training set of ShapeNeRF–Text, and consider such pipelines as natural baselines. Specifically, we use LLaVA (v1.6) [44], and BLIP-2 [39] for images, as well as PointLLM [77] and GPT4Point [58] for colored point clouds. We employ the official codes and pre-trained models released by the respective authors for such evaluations. 111 LLaVA: https://github.com/haotian-liu/LLaVA BLIP-2: https://github.com/salesforce/LAVIS/tree/main/projects/blip2 PointLLM: https://github.com/OpenRobotLab/PointLLM GPT4Point: https://github.com/Pointcept/GPT4Point. . We note that the only official GPT4Point weights available at submission time were those obtained from fine-tuning OPT-2.7B on Cap3D [46]. In the main paper, we present the performance of all methods under the more realistic scenario where NeRFs are treated as the only input data to the assistant. Hence, images and point clouds can only be extracted from NeRFs. Details on the extraction procedure are provided in Sec. A.3. Moreover, in Appendix D, we report the results dealing with the images used to train the NeRF or the original 3D point cloud from ShapeNet, which confirms the methods’ ranking. When rendering an image, a non-obvious design decision for the pipeline is from which vantage point to render it. ShapeNet artificially simplifies this task since all objects have been canonically aligned to a common reference frame, but this may not be the case in a general setting. To show the vantage point’s effect on the assistant’s results, we report results processing a frontal or back view.
5.2 NeRF captioning
In the captioning experiments, we test the assistants’ ability to describe the NeRF content. We prompt them with the NeRF, or the image/cloud extracted from it, followed by the question which has been paired with its ground-truth description, as detailed in section 4.2, e.g. "What’s the content of this NeRF/image/cloud?". We then collect the answers generated by the models and compare them with the ground-truth description according to the selected metrics.
Brief description.
We report the results for the brief description tasks on ShapeNeRF–Text and the HST dataset in Tab. 1 and Tab. 2 respectively. Comparing LLaNA with the baselines described in Sec. 5.1, we appreciate how LLaNA achieves the best performance in most metrics, often by large margins against runner-ups. For instance, for the Sentence-BERT similarity on the ShapeNeRF–Text dataset, LLaNA achieves $68.63$, $7.63$ points more than LLaVA-vicuna13b, even if LLaNA uses a smaller LLM. Results on the HST dataset, which provides ground-truth descriptions validated by humans, are generally lower for all methods. Yet, LLaNA provides again the best performance according to most metrics. The difference in the quality of the brief description provided by LLaNA compared to the baselines is showcased by the qualitative result reported in the first row of Fig. 4, where the description provided by LLaNA is the most accurate.
A clear trend in both tables and qualitative results is that image-based models tend to perform better than models processing point clouds. This is likely due to the larger amount of data used during training of the modality encoder, i.e. millions of images versus hundreds of thousands of shapes, which enhances their generalization ability, as well as the capability of images to capture more details than point clouds at the input resolutions required by image-based MLLMs versus 3D MLLMs. Nonetheless, our method, which operates on NeRFs, benefits from a holistic view of the object and provides the most accurate descriptions. Remarkably, in LLaNA, all the necessary information for this language task can be extracted from a single global embedding obtained by directly processing the NeRF weights. It is also worth pointing out that, while LLaNA directly processes weights and thus is independent by design from spatial resolution, the baselines face a computational overhead growing with the desired resolution due to the necessity of extracting spatial data from NeRF (Sec. A.3). Finally, comparing the results of image-based MLLMs when processing front versus back views, we can see that the vantage point has a non-negligible effect on the performance of such baselines, with SentenceBERT and SimCSE metrics diminishing by about 4 points in all baselines. In a dataset without canonical poses for objects, this would be a relevant limitation that processing NeRF weights seamlessly sidesteps.
Model Modality Sentence-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (Front View) 61.00 61.16 14.30 20.00 23.31
LLaVA-vicuna-13b Image (Back View) 54.35 56.09 21.94 21.67 22.09
LLaVA-vicuna-7b Image (Front View) 59.85 62.35 22.67 23.24 23.35
LLaVA-vicuna-7b Image (Back View) 55.68 58.46 21.97 22.46 22.50
BLIP-2 FlanT5-xxl Image (Front View) 56.13 58.21 5.46 18.69 9.67
BLIP-2 FlanT5-xxl Image (Back View) 52.48 54.05 5.67 18.20 9.50
@@ 5.2 NeRF captioning
Table 1: NeRF captioning - brief description on ShapeNeRF-Text. Baseline results obtained on data extracted from NeRFs. Best results in bold. Runner-up underlined.
Model Modality Sentence-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (Front View) 55.62 55.56 6.56 11.81 14.52
LLaVA-vicuna-13b Image (Back View) 50.00 50.79 9.39 12.76 14.46
LLaVA-vicuna-7b Image (Front View) 54.31 56.28 10.08 14.71 14.53
LLaVA-vicuna-7b Image (Back View) 51.75 52.29 8.13 13.96 14.18
BLIP-2 FlanT5-xxl Image (Front View) 57.11 59.43 8.21 18.02 12.14
BLIP-2 FlanT5-xxl Image (Back View) 54.11 56.37 9.09 17.38 11.79
@@
Table 2: NeRF captioning - brief description on the HST dataset. Baseline results obtained on data extracted from NeRFs. Best results in bold. Runner-up underlined.
Detailed description.
We evaluate the performance for the detailed description tasks on the proposed ShapeNeRF–Text, reporting the results in Tab. 3 and the second row of Fig. 4. For this task, the point-based model PointLLM [77] performs similarly to the image-based one LLaVA [44]. However, we appreciate that LLaNA achieves the best performance in all metrics by large margins. For instance, for the Sentence-BERT metric, LLaNA achieves $77.43$, notably $18.35$ points more than LLaVA-vicuna-13b. These large improvements indicate that, while individual images may be sufficient for brief descriptions, they may lack all the details needed to provide a comprehensive description. Moreover, the dependency of the output quality on the selected vantage point remains strong. Contrarily, the NeRF weights contain detailed and complete information about the object, which is fundamental for more granular description tasks, with the additional advantage of not requiring tuning such hyperparameters. The ability of NeRF to capture holistic information about the object is also shown in the second row of Fig. 4, where only the direct processing of NeRF weights lets LLaNA understand that the object is a TV. PointLLM and LLaVA provide detailed but wrong descriptions, likely because of the need to extract the intermediate discrete representation as a point cloud or an image, losing information. Indeed, in both cases, it is hard even for a human observer to provide the right description from the intermediate modalities shown in the figure.
Model Modality Sentence-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (Front View) 59.08 58.87 23.63 23.55 22.55
LLaVA-vicuna-13b Image (Back View) 50.09 50.33 13.77 21.36 13.18
LLaVA-vicuna-7b Image (Front View) 57.55 57.68 14.99 22.82 14.36
LLaVA-vicuna-7b Image (Back View) 53.11 54.46 14.73 22.47 14.05
BLIP-2 FlanT5-xxl Image (Front View) 41.27 40.69 0.18 7.83 2.60
BLIP-2 FlanT5-xxl Image (Back View) 38.49 37.89 0.19 7.72 2.58
@@
Table 3: NeRF captioning - detailed description on ShapeNeRF–Text. Baseline results obtained on data extracted from NeRFs. Best results in bold. Runner-up underlined.
Figure 4: Qualitative results of NeRF captioning and Q&A. Results on ShapeNeRF-Text. From top to bottom: brief and detailed descriptions, single-round Q&A
Model Modality Sentence-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (Front View) 71.61 70.98 20.19 30.42 32.53
LLaVA-vicuna-13b Image (Back View) 68.25 69.06 20.03 29.84 32.27
LLaVA-vicuna-7b Image (Front View) 71.79 71.96 25.79 34.04 34.86
LLaVA-vicuna-7b Image (Back View) 70.88 70.93 25.17 33.30 34.22
BLIP-2 FlanT5-xxl Image (Front View) 45.20 47.92 11.50 20.16 13.49
BLIP-2 FlanT5-xxl Image (Back View) 45.06 47.66 11.50 19.98 13.44
@@
Table 4: NeRF Q&A - single round conversation on ShapeNeRF–Text. Baseline results obtained on data extracted from NeRFs. Best results in bold. Runner-up underlined.
Figure 5: NeRF multi-round Q&A example.
@@
LLaVA-vicuna-13b Image (Front View) 66.13
LLaVA-vicuna-13b Image (Hard View) 63.90
LLaVA-vicuna-7b Image (Front View) 60.25
LLaVA-vicuna-7b Image (Hard View) 57.00
BLIP-2 FlanT5-xxl Image (Front View) 63.67
BLIP-2 FlanT5-xxl Image (Hard View) 61.47

### v2

@@ 1 Introduction
Large Language Models (LLMs) [1, 70] have revolutionized the field of Natural Language Processing, demonstrating incredible text comprehension and generation capabilities. These results have fostered the development of Multimodal LLMs (MLLMs) [14, 82, 44, 11, 10], which can ingest various modalities such as images, videos, and audio, to generate text describing and reasoning about the content of such modalities. Recently, MLLMs have also been extended to 3D data [58, 78], primarily represented through colored point clouds, yielding remarkable results even in this scenario.
@@
Figure 1: LLaNA. The first Multimodal Large Language Model that understands and reasons on an input NeRF. Our framework directly processes the NeRF weights and performs tasks such as captioning, Q&A, and zero-shot classification of NeRFs.
@@ 2 Related work
Multimodal Large Language Models (MLLMs).
@@
The interaction between NeRF and language has been recently investigated for several practical applications. Many works address the problem of generating geometrically consistent views of objects or scenes described by textual prompts [66, 49, 31, 65, 40, 36, 57]. Other approaches focus on editing the scene represented by a NeRF from text, e.g., by changing the appearance and shape of objects [73, 28, 67, 74, 69, 21, 80, 87], or by inserting/removing objects in the scene [3, 51]. Some techniques investigate new types of radiance fields that predict language features for each spatial location alongside density and color [32, 34]. By distilling knowledge from vision-language models into these models, the neural fields can be queried by textual prompts. LERF [32] extends the original radiance field formulation, considering functions which model density, color and language features at each spatial coordinate. Such language fields are parametrized by a neural network. Unlike all previous methods, Ballerini et al. [5] is the first to utilize the weights of a NeRF’s MLP as an input modality. They aim to learn a mapping between the NeRF and CLIP [59] embedding spaces to perform tasks such as NeRF retrieval from textual or image queries. Differently, our goal is to develop an MLLM capable of reasoning about NeRFs.
@@ 3 Methodology
Neural Radiance Field (NeRF) [50] is a framework that employs coordinate-based neural networks, typically MultiLayer Perceptrons (MLP) and is trained on a collection of images of an object or scene taken from various vantage points. The main application of NeRFs is the task of novel views synthesis, i.e., photorealistic rendering of images from viewpoints unseen at training time. In its base formulation, the MLP is a function of continuous 3D coordinates $\mathbf{p}=(x,y,z)\in\mathbb{R}^{3}$, that yields four-dimensional outputs, $RGB\sigma\in[0,1]^{4}$. This output encodes the $RGB$ color and the volume density $\sigma$ of each 3D location in the scene. The volume density $\sigma$ can be interpreted as the differential probability of a ray terminating at point $\mathbf{p}$. After training, a NeRF can render images from any desired viewpoint at arbitrary resolution by querying it for the values of $RGB$ and $\sigma$ at several points along the ray corresponding to each pixel and applying the volumetric rendering equation [50].
In this work, we realize NeRFs as MLPs composed of $L$ hidden layers, an input layer, and an output layer. An example of MLP with 1 input, 1 output, and 1 hidden layer is shown in Fig. 2 (left). A layer is parameterized by a weight matrix plus a bias vector. More in detail, the hidden layers in our architecture have the same number of input and output neurons, $H$, thus having squared weight matrices $\mathbf{W}_{l}\in\mathbb{R}^{H\times H}$ for $l=1,\dots,L$ and $H$-dimensional biases $\mathbf{b}_{l}\in\mathbb{R}^{H}$. As input $\mathbf{p}$ goes through a 24-frequency encoding [50], the first layer has $\mathbf{W}_{in}\in\mathbb{R}^{144\times H}$ and $\mathbf{b}_{in}\in\mathbb{R}^{H}$. The final one has $\mathbf{W}_{out}\in\mathbb{R}^{H\times 4}$ and $\mathbf{b}_{out}\in\mathbb{R}^{4}$. Refer to Appendix A for more details on NeRFs.
@@
Inspired by the recent development of meta-networks capable of processing neural fields [42, 81], we employ as our meta-encoder architecture nf2vec [81]. It takes as input the weights of a NeRF and yields a global embedding that distills the content of the input NeRF. In particular, the weight matrices and biases of the input NeRF are stacked along the row dimension to form a matrix $\mathbf{M}\in\mathbb{R}^{S\times H}$, where the number of rows $S$ depends on the number of hidden layers $L$, the number of units per hidden layer $H$, and the dimension of the input, which is a $144$-dimensional array obtained by frequency encoding of the 3D coordinates. Before stacking, we pad the output layer weights $\mathbf{W}_{out}$ and biases $\mathbf{b}_{out}$ with zeros to obtain $H$ columns (see Fig. 2, center).
@@
Our framework is trained on the ShapeNeRF–Text dataset, described in detail in Sec. 4. This dataset is organized into a set of prompts from the user and expected ground-truth answers that are used to optimize the original auto-regressive objective of the LLM. For the meta-encoder, we employ the nf2vec encoder pre-trained on ShapeNet released by the authors [81], and we keep it frozen during training. We follow the two-stage training protocol delineated in Liu et al. [44]:
@@
Our model is implemented in PyTorch and trained on 4 NVIDIA A100 with 64GB of VRAM each. Completing both stages requires $\sim$1 day of training.
@@ 4.1 ShapeNeRF–Text dataset
To train and validate our NeRF assistant, we automatically created a dataset of conversations about NeRFs, the ShapeNeRF–Text dataset.
Figure 3: Automatic annotation pipeline. Given a 3D model, $N$ views are rendered and processed by a VLM (LLaVA) to generate view-specific captions. These are aggregated by an LLM (LLaMA) for final descriptions and Q&A.
It features paired NeRFs and language annotations for ShapeNet objects [8], in particular for all the 40K NeRFs available in the nf2vec dataset [61]. We followed the structure defined in PointLLM [78] to create the textual annotations. More in detail, for each object, we generated a brief description, a detailed description, 3 single-round Q&As, and one multi-round Q&A. The brief descriptions are concise captions of the object, taking into account its global structure and appearance. The detailed descriptions are longer sentences that describe all the details of the object. The single-round Q&As consist of a question about the object and the corresponding ground-truth answer. Finally, the multi-round Q&As are longer conversations formed by 3 questions and the relative answers. The automatic data annotation pipeline is inspired by Cap3D [46] and is shown in Fig. 3. First, multiple views of each ShapeNet object have been rendered from different perspectives. Then, each view has been provided as input to LLaVA (LLaVA2-13b) [44] to get a detailed description of the object from that point of view. Afterward, starting from the captions generated by LLaVA, LLaMA 3 (LLaMA3-8B-chat) was used to generate the final ground-truth text data (brief and detailed descriptions, single and multi-round Q&As). Both the frozen LLMs employed to create our benchmark (LLaVA2-13b, LLaMA3-8b-chat) are equipped with safeguards.
When building the ground-truth data, to ensure diversity in the language annotations, each brief and detailed description has been associated with a question randomly sampled from 30 instructions for each kind of description. Such instructions, together with the carefully engineered request prompts for LLaVA and LLaMA, are reported in Sec. C.1.
ShapeNeRF–Text provides 30939, 3846 and 3859 objects for the train, validation and test sets, respectively. Overall, the dataset features 13 object classes, and the train, validation and test splits are obtained by randomly sampling objects within each class, i.e., holding out a fixed percentage of objects per class (80%, 10%, and 10% for the sets, respectively). Sec. C.2 provides more dataset statistics. As quantitatively proven in Sec. C.3 and Sec. D.1, many of the questions belonging to the Q&A set require a holistic 3D understanding of the object, to be answered correctly.
@@ 4.2 Language tasks and metrics
We evaluate NeRF assistants on three different language tasks, given an input NeRF: brief captioning, detailed captioning, and single-round Q&A. We evaluate all tasks on the objects from the ShapeNeRF–Text test set. For brief captioning, we additionally evaluate the methods on the GPT2Shape Human Shape Text (HST) dataset [2], a subset of ShapeNet for which human-curated brief descriptions are publicly available. To generate the dialogues for HST, we randomly pair each of its captions with one of the 30 instructions requesting a brief description, used in ShapeNeRF–Text and reported in Sec. C.1. We employ standard language similarity metrics to evaluate these methods. We compute the cosine similarity between the global embeddings of the generated and ground-truth sentences provided by the pre-trained encoders Sentence-BERT [63] and SimCSE [18]. These metrics based on learned networks are the most effective at measuring the quality of the generated output. We also include standard handcrafted metrics based on n-gram statistics, like BLEU-1 [55], ROUGE-L [43], and METEOR [6].
@@ 5.1 Foundation models as baselines
As our method is the first to investigate language tasks on NeRF, there are no baselines in the literature. However, given a NeRF, a straightforward way to create an assistant for it could be to render an image and use an MLLM capable of ingesting images. Alternatively, we could extract the 3D shape from the NeRF and use one of the recent 3D MLLMs. Hence, in a first set of experiments, we use MLLMs as off-the-shelf foundation models, trained on hundreds of thousands of shapes or millions of images, without performing any fine-tuning on the training set of ShapeNeRF–Text, and consider such pipelines as natural baselines. Specifically, we use LLaVA (v1.6) [44] and BLIP-2 [39] for images, as well as PointLLM [78] and GPT4Point [58] for colored point clouds. Since NeRFs can render arbitrary viewpoints after training, we also include the evaluation of LLaVA [44] in a multi-view scenario. More in detail, we render images from $N$ viewpoints randomly sampled between the set of camera poses used to train each NeRF; then, we concatenate tokens from these N images and fed them into LLaVA alongside text instructions. We set $N$=$3$ because the model cannot process a higher number of images correctly. In addition, we test 3D-LLM [24] to compare its performance to LLaNA. We employ the official code and pre-trained models released by the respective authors for such evaluations 111 LLaVA: https://github.com/haotian-liu/LLaVA BLIP-2: https://github.com/salesforce/LAVIS/tree/main/projects/blip2 PointLLM: https://github.com/OpenRobotLab/PointLLM GPT4Point: https://github.com/Pointcept/GPT4Point 3D-LLM: https://github.com/UMass-Foundation-Model/3D-LLM. We note that the only official GPT4Point weights available at submission time were those obtained from fine-tuning OPT-2.7B on Cap3D [46]. In Tabs. 2, 2, 3, 4, 4 and 5.1, we present the performance of all methods under the more realistic scenario where NeRFs are treated as the only input data to the assistant. Hence, images and point clouds can only be extracted from NeRFs. Details on the extraction procedure are provided in Sec. A.3. As for 3D-LLM, we extract colored 3D meshes from the NeRFs of ShapeNeRF–Text and process such data with the official 3D-LLM code to render images from multiple views and compute both the 2D and 3D features required by the model at inference time. Moreover, in Appendix E, we report the results dealing with the images used to train the NeRF or the original 3D point cloud from ShapeNet, which confirms the methods’ ranking. When rendering an image, a non-obvious design decision for the pipeline is from which vantage point to render it. ShapeNet artificially simplifies this task since all objects have been canonically aligned to a common reference frame, but this may not be the case in a general setting. To show the vantage point’s effect on the assistant’s results, we report results processing a frontal or back view.
Table 1: NeRF brief captioning on ShapeNeRF-Text. Frozen baselines.
Best results are in
bold, runner-up is underlined.
(FV: front-view, BV: back-view, MV: multi-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 61.00 61.16 14.30 20.00 23.31
LLaVA-vicuna-13b Image (BV) 54.35 56.09 21.94 21.67 22.09
LLaVA-vicuna-13b Image (MV) 59.64 61.01 22.84 22.17 23.08
LLaVA-vicuna-7b Image (FV) 59.85 62.35 22.67 23.24 23.35
LLaVA-vicuna-7b Image (BV) 55.68 58.46 21.97 22.46 22.50
BLIP-2 FlanT5-xxl Image (FV) 56.13 58.21 5.46 18.69 9.67
BLIP-2 FlanT5-xxl Image (BV) 52.48 54.05 5.67 18.20 9.50
@@
Table 2: NeRF brief captioning on the HST dataset. Frozen baselines.
Best results are in
bold, runner-up is underlined.
(FV: front-view, BV: back-view, MV: multi-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 55.62 55.56 6.56 11.81 14.52
LLaVA-vicuna-13b Image (BV) 50.00 50.79 9.39 12.76 14.46
LLaVA-vicuna-13b Image (MV) 54.25 55.56 9.78 14.13 14.99
LLaVA-vicuna-7b Image (FV) 54.31 56.28 10.08 14.71 14.53
LLaVA-vicuna-7b Image (BV) 51.75 52.29 8.13 13.96 14.18
BLIP-2 FlanT5-xxl Image (FV) 57.11 59.43 8.21 18.02 12.14
BLIP-2 FlanT5-xxl Image (BV) 54.11 56.37 9.09 17.38 11.79
@@
Table 3: NeRF detailed captioning on ShapeNeRF-Text. Frozen baselines.
Best results are in bold, runner-up is underlined. (FV: front-view, BV: back-view, MV: multi-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 59.08 58.87 23.63 23.55 22.55
LLaVA-vicuna-13b Image (BV) 50.09 50.33 13.77 21.36 13.18
LLaVA-vicuna-13b Image (MV) 60.21 59.51 15.07 32.16 14.64
LLaVA-vicuna-7b Image (FV) 57.55 57.68 14.99 22.82 14.36
LLaVA-vicuna-7b Image (BV) 53.11 54.46 14.73 22.47 14.05
BLIP-2 FlanT5-xxl Image (FV) 41.27 40.69 0.18 7.83 2.60
BLIP-2 FlanT5-xxl Image (BV) 38.49 37.89 0.19 7.72 2.58
@@
Table 4: NeRF single-round Q&A on ShapeNeRF-Text. Frozen baselines.
Best results are in bold, runner-up is underlined. (FV: front-view, BV: back-view, MV: multi-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 71.61 70.98 20.19 30.42 32.53
LLaVA-vicuna-13b Image (BV) 68.25 69.06 20.03 29.84 32.27
LLaVA-vicuna-13b Image (MV) 71.84 71.16 20.04 30.20 33.46
LLaVA-vicuna-7b Image (FV) 71.79 71.96 25.79 34.04 34.86
LLaVA-vicuna-7b Image (BV) 70.88 70.93 25.17 33.30 34.22
BLIP-2 FlanT5-xxl Image (FV) 45.20 47.92 11.50 20.16 13.49
BLIP-2 FlanT5-xxl Image (BV) 45.06 47.66 11.50 19.98 13.44
@@
Figure 4: Qualitative results of NeRF captioning and Q&A. Results on ShapeNeRF–Text. From top to bottom: brief and detailed descriptions, single-round Q&A
Table 5: Zero-Shot NeRF Classification. Frozen baselines.
Best results are in bold, runner-up is underlined.
(FV: front-view, BV: back-view, MV: multi-view)
@@
LLaVA-vicuna-13b Image (FV) 66.13
LLaVA-vicuna-13b Image (BV) 63.90
LLaVA-vicuna-13b Image (MV) 73.45
LLaVA-vicuna-7b Image (FV) 60.25
LLaVA-vicuna-7b Image (BV) 57.00
BLIP-2 FlanT5-xxl Image (FV) 63.67
BLIP-2 FlanT5-xxl Image (BV) 61.47